# A Sharp condition on global wellposedness of Chern-Simon-Shrödinger equation

### v1

@@ Abstract.
In this work, we derive a sharp condition on the mass of the initial data for the global existence of the Chern-Simon-Shrödinger equation. As a corollary, we prove that if the strength of interaction is less than the Bogomolny bound, then, for a large enough mass of initial data, there exists a globally defined solution. On the other hand, for the interactions which are above the Bogomolny bound, the critical mass condition on the initial data for the global existence depends on the strength of the self-interacting field. Then, we show that the states with the initial critical mass and zero energy are standing wave solutions and globally well-posed. Moreover, they are static if the self-interacting field is large enough and non-static for small self-interacting field.
@@
Golbal wellposedness, Chern-Simon-Shrödinger equation, Standing and static wave solutions
@@ 1. Introduction
The Chern-Simon-Shrödinger equation has its origin in quantum mechanics as a model for anyons and fractional quantum Hall effect; see [33, 14, 18, 32, 6, 7, 16, 15, 17, 5] and references therein. It has also been studied in mathematics as a challenging model which contains a non-local and non-homogenous self-generating magnetic energy; see [2, 11, 3, 12, 27, 30, 26, 4, 24, 9, 22, 20] and references therein. The model can be described as follows: Let $\beta\in\mathbb{R}_{+},\gamma\in\mathbb{R}$, and $\phi\in H^{1}(\mathbb{R}^{2})=H^{1}(\mathbb{R}^{2};\mathbb{C})$. For every $f\in L^{1}(\mathbb{R}^{2})$ and $\mathbf{F}\in L^{1}(\mathbb{R}^{2};\mathbb{C}^{2})$, define, in the principal value sense, the self-generating gauges
@@
The parameter $\beta$ shows the strength of the self-interacting magnetic field and $\gamma$ shows the strength of the interaction between the particles, where $\gamma>0$ for attractive particles and $\gamma<0$ for repulsive particles. Now, for $T>0,$ we say that $\psi:[0,T]\times\mathbb{R}^{2}\to\mathbb{R}$ satisfies the Chern-Simon-Shrödinger equation if $\psi\in L^{\infty}([0,T],H^{1}(\mathbb{R}^{2}))\cap C([0,T],L^{2}(\mathbb{R}^{2}))$ and
@@
One of the strange properties of (1.2) is the that if $\beta>0,$ $\|\psi_{0}\|^{2}_{L^{2}}\geq\frac{2}{\beta},$ and $\gamma<2\pi\beta$, then there exists always a global solution for (1.1) with the initial data $\psi_{0}$. Here, we are using $\gamma_{\ast}(\theta)=2\pi\theta$ if $\theta\geq 2$; see [1, Thm. 2]. This is unlike the classical nonlinear Shrödinger equation, where the mass of the initial data should be small to have a global solution; see [31].
@@
To mention some previous works on the global existence, the case of $\beta=0$ and $\gamma=1$, the classical nonlinear Shrödinger equation, has been studied in the fundamental work of Weinstein [31]. He proved in [31, Thm. A,Thm. 4.1] that a global solution to (1.1) with the initial data $\psi_{0}\in H^{1}(\mathbb{R}^{2})$ exists if
@@
and this bound is sharp, where $\gamma_{\ast}(0)$ is the Ladyzhenskaya–Gagliardo–Nirenberg interpolation constant. The global well-posedness for the case of general $\beta$ has been first studied in [2]. The authors considered the lower-bound $\gamma_{\ast}(0)$ for $\gamma_{\ast}\left(\beta\|\psi_{0}\|_{L^{2}}^{2}\right)$ in (1.2). Afterwards, there were several attempts to sharpen this result. The case of radial symmetric solutions and their global behavior was studied in [3]. In [30], the global well-posedness has been proved with initial data which has small $H^{2}$ and a special weighted Sobolev norm. In [26, 9], the global well-posedness for the equivariant case has been studied.
However, the literature is lacking the sharp condition for the global well-posedness of solutions to (1.1). To derive the sharp constant, we use that the interpolation constant $\gamma_{\ast}$ is continuous, the function $\frac{\gamma_{\ast}(\theta)}{\theta}$ is decreasing for $\theta>0,$ and a well-known argument in [31] derives a uniform bound on the $H^{1}$-norm of an approximating solutions, constructed in [2], for (1.1). Then, by a convergence argument, we obtain a global solution. We remark that there is a subtle difference between the way we pass to the limit in compare to [2], namely we do not use any estimate on time derivative of the gauge term $\mathbf{A}[|\psi|^{2}]$. Instead, we use a local convergence property for the gauge terms; see Lemma 3.3. Finally, the other part of the theorem follows as the solutions with negative energy blows up at finite time; see [2, Thm 3.1].
@@
Let $\beta,\gamma\in\mathbb{R}$, and $\psi_{0}\in H^{1}(\mathbb{R}^{2})\setminus 0$ satisfy
@@
Let $\beta,\gamma\in\mathbb{R},T>0$, and $\psi_{0}\in H^{1}(\mathbb{R}^{2}).$ Then, if $\int_{\mathbb{R}^{2}}|\mathbf{x}|^{2}|\psi_{0}(\mathbf{x})|^{2}\,{\rm d}\mathbf{x}<\infty$ and a solution $\psi\in C([0,T];H^{2}(\mathbb{R}^{2}))$ of (1.1) with the initial data $\psi_{0}$ is standing wave solution, i.e.,
@@
In Theorem 1.2, it is crucial to start with initial data $\psi_{0}$ which has zero energy. Otherwise, by assuming only $\gamma=\frac{\gamma_{\ast}\left(\beta\|\psi_{0}\|_{L^{2}}^{2}\right)}{\|\psi_{0}\|_{L^{2}}^{2}}$, by using conformal transformations, we may derive solutions which blow up in finite time; see [11, 21, 19, 20].
@@ 3. Preliminary results
Let $u\in H^{1}(\mathbb{R}^{2})$. Then, the following inequalities hold.
@@
    $|\nabla|u|(\mathbf{x})|\leq\left|(\nabla+{\rm i}\beta\mathbf{A}\left[|u|^{2}\right])u(\mathbf{x})\right|,$ (3.1)
@@
    $\int_{\mathbb{R}^{2}}\left|\mathbf{A}\left[|u|^{2}\right]u\right|^{2}\leq C_{\rm H}\left(\int_{\mathbb{R}^{2}}|u|^{2}\right)^{2}\left(\int_{\mathbb{R}^{2}}\bigl{|}\nabla|u|\bigr{|}^{2}\right),$ (3.2)
@@
    $C(p)\int_{\mathbb{R}^{2}}|u|^{p}\leq\left(\int_{\mathbb{R}^{2}}|u|^{2}\right)\left(\int_{\mathbb{R}^{2}}|\nabla u|^{2}\right)^{\frac{p-2}{2}},$ (3.3)
@@
Let $(\lambda,\beta,\gamma)\in\mathbb{R}_{+}^{3}$, $T>0,$ and $\psi_{0}\in H^{2}(\mathbb{R}^{2})$. Assume that $\psi\in C([0,T],H^{2}(\mathbb{R}^{2}))$ is a weak solution of
@@
The proof of (3.5) follows the same argument as [2, Prop 2.1], so we avoid repetition. Finally, to demonstrate (3.4), we can multiply the equation for $\psi$ with $\psi$ and integrate the imaginary part, as mentioned in [2, Page 245]. ∎
@@ 4. Proof of the theorem
$\psi^{\varepsilon}_{0}=(1-\varepsilon^{\frac{1}{4}}\Delta)^{-1}\psi_{0}.$
By the same argument as [2, Lem. 2.2], there exists a unique $\psi^{\varepsilon}\in C(\mathbb{R}_{+},H^{2}(\mathbb{R}^{2}))$ solution of
@@
in +×$\mathbb{R}_{+}\times\mathbb{R}^{2}$, with the initial value $\psi_{0}^{\varepsilon}.$ Note that, by a simple application of the Fourier transform, we have
$\lim_{\varepsilon\to 0}\|\psi_{0}^{\varepsilon}-\psi_{0}\|_{H^{1}(\mathbb{R}^{2})}=0.$
Moreover, by Lemma 3.2, the assumption (1.2), and the continuity of $\gamma_{\ast}$; see [1, Prop. 3.26], for $\delta>0,\varepsilon_{0}>0$ small enough, we have
$\gamma(1+\delta)<\frac{\gamma_{\ast}\left(\beta(1+4\sqrt{\varepsilon})\|\psi_{0}\|_{L^{2}}^{2}\right)}{(1+4\sqrt{\varepsilon})\|\psi_{0}\|_{L^{2}}^{2}},$
for every $0<\varepsilon<\varepsilon_{0}.$ In the rest of the argument, we consider $0<\varepsilon<\max\left(\varepsilon_{0},1\right)$ such that (LABEL:eq:choiceofepsilon) holds. Now, we derive a global bound on $\|\psi^{\varepsilon}(t,\cdot)\|_{H^{1}(\mathbb{R}^{2})}$. To prove the bound, we first use the conservation of the mass and energy in Proposition 3.1 to imply that
@@
for every $t\in\mathbb{R}_{+}$. Then, by (LABEL:eq:choiceofepsilon), we obtain
@@
$=\|\psi_{0}\|^{2}_{L^{2}}+\sqrt{\varepsilon}\left\|\psi_{0}-(1-\varepsilon^{\frac{1}{4}}\Delta)^{-1}\psi_{0}\right\|^{2}_{L^{2}}\leq(1+4\sqrt{\varepsilon})\|\psi_{0}\|^{2}_{L^{2}},$
@@
$\frac{\gamma_{\ast}\left(\beta\|\psi^{\varepsilon}(t,\cdot)\|^{2}_{L^{2}}\right)}{\|\psi^{\varepsilon}(t,\cdot)\|^{2}_{L^{2}}}\geq\frac{\gamma_{\ast}\left(\beta(1+4\sqrt{\varepsilon})\|\psi_{0}\|^{2}_{L^{2}}\right)}{(1+4\sqrt{\varepsilon})\|\psi_{0}\|^{2}_{L^{2}}}.$
for every $t\in\mathbb{R}_{+}.$ Therefore, by (LABEL:eq:choiceofepsilon), we derive that
@@
for every $t\in\mathbb{R}_{+}.$ Then, by using (4.3) and (4.5), we obtain
@@
where we used [1, Lem. 3.3] in the last inequality. Therefore, by (LABEL:eq:choiceofepsilon), (4.4), and Gagliardo-Nirenberg interpolation inequality (3.3), $\|\psi^{\varepsilon}(t,\cdot)\|_{H^{1}}$ is uniformly bounded and
@@
for all $t\geq 0$, where $C_{1}$ is a constant depending on $\beta,\gamma,\delta,\|\psi_{0}\|_{H^{1}}$. Now, we aim to prove that every term on the right-hand-side of (LABEL:eq:approximatingsolution) is uniformly bounded in
@@
To do so, we need to show that $\|\mathbf{A}[|\psi^{\varepsilon}(t,\cdot)|^{2}]\|_{L^{2}},\|\mathbf{A}[|\psi^{\varepsilon}(t,\cdot)|^{2}]\|_{L^{4}},\|\mathbf{A}\ast[\textbf{j}_{\beta}[\psi^{\varepsilon}(t,\cdot)]]\|_{L^{4}}$ are uniformly bounded for $t\in\mathbb{R}_{+}.$
@@
$\leq C_{2}\,C_{3}\|\psi^{\varepsilon}(t,\cdot)\|^{\frac{1}{2}}_{L^{2}}\|\nabla\psi^{\varepsilon}(t,\cdot)\|^{\frac{1}{3}}_{L^{2}}\leq C_{1}^{\frac{1}{3}}\,C_{2}\,C_{3}.$
@@
$\leq C_{2}\left(C_{1}\,C_{4}+\beta C_{1}^{\frac{4}{3}}\,C_{2}^{2}\,C_{3}C_{4}\right).\,$
Here, we used Proposition 3.3 on the first inequality, Hölder’s inequality on the third inequality, (4.7) and Gagliardo-Nirenberg interpolation inequality (3.3) on the fourth inequality, and (4.6) on the last inequality. In conclusion, by Lemma 3.2 $(iii)$ and (LABEL:eq:approximatingsolution), we derive that $\partial_{t}\psi^{\varepsilon}$ is uniformly bounded in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+},H^{-1}(\mathbb{R}^{2})),$ and $\psi^{\varepsilon}$ is uniformly bounded in $L^{\infty}(\mathbb{R}_{+};H^{1}(\mathbb{R}^{2}))$. Then, by a classical theorem of Lions, see [25], we conclude that, up to a subsequence, $\psi^{\varepsilon}$ converges to $\psi\in L^{\infty}(\mathbb{R}_{+},H^{1}(\mathbb{R}^{2}))\cap C(\mathbb{R}_{+},L^{2}(\mathbb{R}^{2}))$ weakly in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+},H^{1}(\mathbb{R}^{2}))$, pointwise a.e. in +×$\mathbb{R}_{+}\times\mathbb{R}^{2}$, and strongly in the space $L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$ for every $1\leq q<\infty$.
Now, we claim that $\psi$ satisfies (1.1) weakly in +×$\mathbb{R}_{+}\times\mathbb{R}^{2}$ with the initial data $\psi_{0}.$ To prove the claim, we use the convergence of $\psi^{\varepsilon}_{0}$ to $\psi_{0}$ to derive that $\psi(0,\cdot)=\psi_{0}(\cdot).$ Hence, it is sufficient to prove that the right hand side of (LABEL:eq:approximatingsolution) converges to the right hand side of (1.1) in the distribution sense. Let $T>0$ and $\varphi\in C([0,T];C^{\infty}_{c}(\mathbb{R}^{2}))$ be a test function. Since, by (4.4), $\sup_{t\in\mathbb{R}_{+}}\left\||\psi_{\varepsilon}(t,\cdot)|^{2}\right\|_{L^{1}}$ is uniformly bounded and $\psi_{\varepsilon}$ converges to $\psi$ in $L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$ for every $1\leq q<\infty$, by Lemma 3.3, we derive that $\mathbf{A}[|\psi_{\varepsilon}|^{2}]$ converges to $\mathbf{A}[|\psi|^{2}]$ in $L^{p}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$ for every $0<p<\infty$. Then,
@@
To show the converges of the term $(\mathbf{A}\ast\textbf{j}_{\beta}[[\psi^{\varepsilon}]])\psi^{\varepsilon}$ in the distribution, we need more work. First, we note that $\textbf{j}_{\beta}[\psi^{\varepsilon}]=\mathbf{J}[\psi^{\varepsilon}]+\beta\mathbf{A}[|\psi^{\varepsilon}|^{2}]|\psi^{\varepsilon}|^{2}.$ Then, it is enough to prove the convergence of $(\mathbf{A}\ast[\mathbf{J}[\psi^{\varepsilon}]])\psi^{\varepsilon}$ and $(\mathbf{A}\ast[\mathbf{A}[|\psi^{\varepsilon}|^{2}]|\psi^{\varepsilon}|^{2}])\psi^{\varepsilon}$ separately. For the second one, we use Lemma 3.3. To prove the conditions of Lemma 3.3, we consider
@@
where Hölder’s inequality is used on the first inequality, (4.7) and Gagliardo-Nirenberg’s inequality (3.3) are used on the second inequality, and (4.6) is used on the last inequality. Moreover, since $\mathbf{A}[|\psi^{\varepsilon}|^{2}],\psi^{\varepsilon}$ converges to $\mathbf{A}[|\psi|^{2}],\psi$ in $L^{p}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2};\mathbb{R}^{2}),L^{p}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$, respectively, for every $1\leq p<\infty$, we derive that $\mathbf{A}[|\psi^{\varepsilon}|^{2}]|\psi^{\varepsilon}|^{2}$ converges to $\mathbf{A}[|\psi|^{2}]|\psi|^{2}$ in $L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2};\mathbb{R}^{2})$ for every $1\leq q<\infty$. Hence, by Lemma 3.3, it is implied that $\mathbf{A}\ast[\mathbf{A}[|\psi^{\varepsilon}|^{2}]|\psi^{\varepsilon}|^{2}]$ converges to $\mathbf{A}\ast[\mathbf{A}[|\psi|^{2}]|\psi|^{2}]$ in $L^{p}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$ for every $0<p<\infty$.

### v2

# A Sharp condition on global wellposedness of Chern-Simons-Schrödinger equation

@@ Abstract.
In this work, we derive a sharp condition on the mass of the initial data for the global existence of the Chern-Simons-Schrödinger equation. As a corollary, we prove that if the strength of interaction is less than the Bogomolny bound, then, for a large enough mass of initial data, there exists a globally defined solution. On the other hand, for the interactions which are above the Bogomolny bound, the critical mass condition on the initial data for the global existence depends on the strength of the self-interacting field. Then, we show that the states with the initial critical mass and zero energy are standing wave solutions and globally well-posed. Moreover, they are static if the self-interacting field is large enough and non-static for small self-interacting field.
@@
Golbal wellposedness, Chern-Simons-Schrödinger equation, Standing and static wave solutions
@@ 1. Introduction
The Chern-Simons-Schrödinger equation has its origin in quantum mechanics as a model for anyons and fractional quantum Hall effect; see [33, 14, 18, 32, 6, 7, 16, 15, 17, 5] and references therein. It has also been studied in mathematics as a challenging model which contains a non-local and non-homogenous self-generating magnetic energy; see [2, 11, 3, 12, 27, 30, 26, 4, 24, 9, 22, 20] and references therein. The model can be described as follows: Let $\beta\in\mathbb{R}_{+},\gamma\in\mathbb{R}$, and $\phi\in H^{1}(\mathbb{R}^{2})=H^{1}(\mathbb{R}^{2};\mathbb{C})$. For every $f\in L^{1}(\mathbb{R}^{2})$ and $\mathbf{F}\in L^{1}(\mathbb{R}^{2};\mathbb{C}^{2})$, define, in the principal value sense, the self-generating gauges
@@
The parameter $\beta$ shows the strength of the self-interacting magnetic field and $\gamma$ shows the strength of the interaction between the particles, where $\gamma>0$ for attractive particles and $\gamma<0$ for repulsive particles. Now, for $T>0,$ we say that $\psi:[0,T]\times\mathbb{R}^{2}\to\mathbb{C}$ satisfies the Chern-Simons-Schrödinger equation if $\psi\in L^{\infty}([0,T],H^{1}(\mathbb{R}^{2}))\cap C([0,T],L^{2}(\mathbb{R}^{2}))$ and
@@
One of the strange properties of (1.2) is the that if $\beta>0,$ $\|\psi_{0}\|^{2}_{L^{2}}\geq\frac{2}{\beta},$ and $\gamma<2\pi\beta$, then there exists always a global solution for (1.1) with the initial data $\psi_{0}$. Here, we are using $\gamma_{\ast}(\theta)=2\pi\theta$ if $\theta\geq 2$; see [1, Thm. 2]. This is unlike the classical nonlinear Schrödinger equation, where the mass of the initial data should be small to have a global solution; see [31].
@@
To mention some previous works on the global existence, the case of $\beta=0$ and $\gamma=1$, the classical nonlinear Schrödinger equation, has been studied in the fundamental work of Weinstein [31]. He proved in [31, Thm. A,Thm. 4.1] that a global solution to (1.1) with the initial data $\psi_{0}\in H^{1}(\mathbb{R}^{2})$ exists if
@@
and this bound is sharp, where $\gamma_{\ast}(0)$ is the Ladyzhenskaya–Gagliardo–Nirenberg interpolation constant. The global well-posedness for the case of general $\beta$ has been first studied in [2]. The authors considered the lower-bound $\gamma_{\ast}(0)$ for $\gamma_{\ast}\left(\beta\|\psi_{0}\|_{L^{2}}^{2}\right)$ in (1.2). Afterwards, there were several attempts to sharpen this result. The case of radially symmetric solutions and their global behavior was studied in [3]. In [30], the global well-posedness has been proven with initial data which has small $H^{2}$ and a special weighted Sobolev norm. In [26, 9], the global well-posedness for the equivariant case has been studied.
However, the literature is lacking the sharp condition for the global well-posedness of solutions to (1.1). To derive the sharp constant, we use that the interpolation constant $\gamma_{\ast}$ is continuous, the function $\frac{\gamma_{\ast}(\theta)}{\theta}$ is decreasing for $\theta>0,$ and a well-known argument in [31] derives a uniform bound on the $H^{1}$-norm of an approximating solutions, constructed in [2], for (1.1). Then, by a convergence argument, we obtain a global solution. We remark that there is a subtle difference between the way we pass to the limit in compare to [2], namely we do not use any estimates on the time derivative of the gauge term $\mathbf{A}[|\psi|^{2}]$. Instead, we use a local convergence property for the gauge terms; see Lemma 3.3. Finally, the other part of the theorem follows as the solutions with negative energy blows up at finite time; see [2, Thm 3.1].
@@
Let $\beta\in\mathbb{R}_{+},\gamma\in\mathbb{R}$, and $\psi_{0}\in H^{1}(\mathbb{R}^{2})\setminus 0$ satisfy
@@
Let $\beta\in\mathbb{R}_{+},\gamma\in\mathbb{R},T>0$, and $\psi_{0}\in H^{1}(\mathbb{R}^{2}).$ Then, if $\int_{\mathbb{R}^{2}}|\mathbf{x}|^{2}|\psi_{0}(\mathbf{x})|^{2}\,{\rm d}\mathbf{x}<\infty$ and a solution $\psi\in C([0,T];H^{2}(\mathbb{R}^{2}))$ of (1.1) with the initial data $\psi_{0}$ is standing wave solution, i.e.,
@@
In Theorem 1.2, it is crucial to start with an initial data $\psi_{0}$ which has zero energy. Otherwise, by assuming only $\gamma=\frac{\gamma_{\ast}\left(\beta\|\psi_{0}\|_{L^{2}}^{2}\right)}{\|\psi_{0}\|_{L^{2}}^{2}}$, by using conformal transformations, we may derive solutions which blow up in finite time; see [11, 21, 19, 20].
@@ 3. Preliminary results
Let $f\in H^{1}(\mathbb{R}^{2})$. Then, the following inequalities hold.
@@
    $|\nabla|f|(\mathbf{x})|\leq\left|(\nabla+{\rm i}\beta\mathbf{A}\left[|f|^{2}\right])f(\mathbf{x})\right|,$ (3.1)
@@
    $\int_{\mathbb{R}^{2}}\left|\mathbf{A}\left[|f|^{2}\right]f\right|^{2}\leq C_{\rm H}\left(\int_{\mathbb{R}^{2}}|f|^{2}\right)^{2}\left(\int_{\mathbb{R}^{2}}\bigl{|}\nabla|f|\bigr{|}^{2}\right),$ (3.2)
@@
    $C(p)\int_{\mathbb{R}^{2}}|f|^{p}\leq\left(\int_{\mathbb{R}^{2}}|f|^{2}\right)\left(\int_{\mathbb{R}^{2}}|\nabla f|^{2}\right)^{\frac{p-2}{2}},$ (3.3)
@@
Let $(\lambda,\beta,\gamma)\in\mathbb{R}_{+}\times\mathbb{R}^{2}$, $T>0,$ and $\psi_{0}\in H^{2}(\mathbb{R}^{2})$. Assume that $\psi\in C([0,T],H^{2}(\mathbb{R}^{2}))$ is a weak solution of
@@
The proof of (3.5) follows the same argument as [2, Prop 2.1], so we avoid repetition. Finally, to demonstrate (3.4), we can multiply the equation for $\psi$ with $\psi^{*}$ and integrate the imaginary part, as mentioned in [2, Page 245]. ∎
@@ 4. Proof of the theorem
$\psi^{\varepsilon}_{0}:=(1-\varepsilon^{\frac{1}{4}}\Delta)^{-1}\psi_{0}.$
Then, by Lemma 3.2,
Moreover, by the same argument as [2, Lem. 2.2], there exists a unique $\psi^{\varepsilon}\in C(\mathbb{R}_{+},H^{2}(\mathbb{R}^{2}))$ solution of
@@
in +×$\mathbb{R}_{+}\times\mathbb{R}^{2}$, with the initial value $\psi_{0}^{\varepsilon}.$ Then, by the assumption (1.2) and the continuity of $\gamma_{\ast}$; see [1, Prop. 3.26], for $\delta>0,\varepsilon_{0}>0$ small enough, we have
$\gamma(1+\delta)<\frac{\gamma_{\ast}\left(\beta(1+\sqrt{\varepsilon})\|\psi_{0}\|_{L^{2}}^{2}\right)}{(1+\sqrt{\varepsilon})\|\psi_{0}\|_{L^{2}}^{2}},$ (4.3)
for every $0<\varepsilon<\varepsilon_{0}.$ In the rest of the argument, we consider $0<\varepsilon<\max\left(\varepsilon_{0},1\right)$ such that (4.3) holds.
Now, we derive a global bound on $\|\psi^{\varepsilon}(t,\cdot)\|_{H^{1}(\mathbb{R}^{2})}$. To prove the bound, we first use the conservation of the mass and energy in Proposition 3.1 to imply that
@@
for every $t\in\mathbb{R}_{+}$. Then, by applying the Fourier transform again, we obtain
@@
$\leq\|\psi_{0}\|^{2}_{L^{2}}+\sqrt{\varepsilon}\left\|\psi_{0}\right\|^{2}_{L^{2}}=(1+\sqrt{\varepsilon})\|\psi_{0}\|^{2}_{L^{2}},$
@@
$\frac{\gamma_{\ast}\left(\beta\|\psi^{\varepsilon}(t,\cdot)\|^{2}_{L^{2}}\right)}{\|\psi^{\varepsilon}(t,\cdot)\|^{2}_{L^{2}}}\geq\frac{\gamma_{\ast}\left(\beta(1+\sqrt{\varepsilon})\|\psi_{0}\|^{2}_{L^{2}}\right)}{(1+\sqrt{\varepsilon})\|\psi_{0}\|^{2}_{L^{2}}}.$
for every $t\in\mathbb{R}_{+}.$ Therefore, by (4.3), it is implied that
@@
for every $t\in\mathbb{R}_{+}.$ Then, by (4.4) and (4.6), we obtain
@@
where we used [1, Lem. 3.3] in the last inequality. Therefore, by (4.1), (4.5), and the Gagliardo-Nirenberg interpolation inequality (3.3), $\|\psi^{\varepsilon}(t,\cdot)\|_{H^{1}}$ is uniformly bounded and
@@
for all $t\in\mathbb{R}_{+}$, where $C_{1}$ is a constant depending on $\beta,\gamma,\delta,\|\psi_{0}\|_{H^{1}}$. Now, we aim to prove that every term on the right-hand-side of (LABEL:eq:approximatingsolution) is uniformly bounded in
@@
To do so, we need to show that $\left\|\mathbf{A}[|\psi^{\varepsilon}(t,\cdot)|^{2}]\right\|_{L^{2}},\left\|\mathbf{A}[|\psi^{\varepsilon}(t,\cdot)|^{2}]\right\|_{L^{4}},\left\|\mathbf{A}\ast[\textbf{j}_{\beta}[\psi^{\varepsilon}(t,\cdot)]]\right\|_{L^{4}}$ are uniformly bounded for $t\in\mathbb{R}_{+}.$
@@
$\leq C_{2}\,C_{3}\|\psi^{\varepsilon}(t,\cdot)\|^{\frac{1}{2}}_{L^{2}}\|\nabla\psi^{\varepsilon}(t,\cdot)\|^{\frac{1}{3}}_{L^{2}}\leq C_{1}^{\frac{5}{12}}\,C_{2}\,C_{3}.$
@@
$\leq C_{1}\,C_{2}\,C_{4}+\beta C_{1}^{\frac{4}{3}}\,C_{2}^{2}\,C_{3}C_{4}.\,$
Here, we used Proposition 3.3 on the first inequality, Hölder’s inequality on the third inequality, (4.8) and the Gagliardo-Nirenberg interpolation inequality (3.3) on the fourth inequality, and (4.7) on the last inequality. In conclusion, by Lemma 3.2 $(iii)$ and (LABEL:eq:approximatingsolution), we derive that $\partial_{t}\psi^{\varepsilon}$ is uniformly bounded in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+},H^{-1}(\mathbb{R}^{2})),$ and $\psi^{\varepsilon}$ is uniformly bounded in $L^{\infty}(\mathbb{R}_{+};H^{1}(\mathbb{R}^{2}))$. Then, by a classical theorem of Lions, see [25], we conclude that, up to a subsequence, $\psi^{\varepsilon}$ converges to $\psi\in L^{\infty}(\mathbb{R}_{+},H^{1}(\mathbb{R}^{2}))\cap C(\mathbb{R}_{+},L^{2}(\mathbb{R}^{2}))$ weakly in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+},H^{1}(\mathbb{R}^{2}))$, pointwise a.e. in +×$\mathbb{R}_{+}\times\mathbb{R}^{2}$, and strongly in the space $L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$ for every $1\leq q<\infty$.
Now, we claim that $\psi$ satisfies (1.1) weakly in +×$\mathbb{R}_{+}\times\mathbb{R}^{2}$ with the initial data $\psi_{0}.$ To prove the claim, by using the Fourier transform, we have the convergence of $\psi^{\varepsilon}_{0}$ to $\psi_{0}$ in $H^{1}(\mathbb{R}^{2})$, which derives that $\psi(0,\cdot)=\psi_{0}(\cdot).$ Hence, it is sufficient to prove that the right hand side of (LABEL:eq:approximatingsolution) converges to the right hand side of (1.1) in the distribution sense. Let $T>0$ and $\varphi\in C([0,T];C^{\infty}_{c}(\mathbb{R}^{2}))$ be a test function. Since, by (4.5), $\sup_{t\in\mathbb{R}_{+}}\left\||\psi_{\varepsilon}(t,\cdot)|^{2}\right\|_{L^{1}}$ is uniformly bounded and $\psi_{\varepsilon}$ converges to $\psi$ in $L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$ for every $1\leq q<\infty$, by Lemma 3.3, we derive that $\mathbf{A}[|\psi_{\varepsilon}|^{2}]$ converges to $\mathbf{A}[|\psi|^{2}]$ in $L^{p}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$ for every $0<p<\infty$. Then,
@@
To show the converges of the term $(\mathbf{A}\ast\textbf{j}_{\beta}[[\psi^{\varepsilon}]])\psi^{\varepsilon}$ in the distribution, we need more work. First, we note that $\textbf{j}_{\beta}[\psi^{\varepsilon}]=\mathbf{J}[\psi^{\varepsilon}]+\beta\mathbf{A}[|\psi^{\varepsilon}|^{2}]|\psi^{\varepsilon}|^{2}.$ Hence, it is enough to prove the convergence of $(\mathbf{A}\ast[\mathbf{J}[\psi^{\varepsilon}]])\psi^{\varepsilon}$ and $(\mathbf{A}\ast[\mathbf{A}[|\psi^{\varepsilon}|^{2}]|\psi^{\varepsilon}|^{2}])\psi^{\varepsilon}$ separately. For the second one, we use Lemma 3.3. To prove the conditions of Lemma 3.3, we consider
@@
where Hölder’s inequality is used on the first inequality, (4.8) and the Gagliardo-Nirenberg’s inequality (3.3) are used on the second inequality, and (4.7) is used on the last inequality. Moreover, since $\mathbf{A}[|\psi^{\varepsilon}|^{2}],\psi^{\varepsilon}$ converges to $\mathbf{A}[|\psi|^{2}],\psi$ in $L^{p}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2};\mathbb{R}^{2}),L^{p}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$, respectively, for every $1\leq p<\infty$, we derive that $\mathbf{A}[|\psi^{\varepsilon}|^{2}]|\psi^{\varepsilon}|^{2}$ converges to $\mathbf{A}[|\psi|^{2}]|\psi|^{2}$ in $L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2};\mathbb{R}^{2})$ for every $1\leq q<\infty$. Hence, by Lemma 3.3, it is implied that $\mathbf{A}\ast[\mathbf{A}[|\psi^{\varepsilon}|^{2}]|\psi^{\varepsilon}|^{2}]$ converges to $\mathbf{A}\ast[\mathbf{A}[|\psi|^{2}]|\psi|^{2}]$ in $L^{p}_{\mathrm{loc}}(\mathbb{R}_{+}\times\mathbb{R}^{2})$ for every $0<p<\infty$.